# Investigating Feature and Model Importance in Android Malware Detection: An Implemented Survey and Experimental Comparison of ML-Based Methods

### v3

@@ 2.3. Evaluation Metrics
(2) - True Positive Rate (TPR):
@@
(4) - True Negative Rate (TNR):
@@
In general, it is necessary to use a range of metrics in order to give a full picture of the performance of different ML models (Lones, 2021). With this in mind, we report the accuracy, precision, F1-score, TPR and TNR for each model. TPR and TNR, in particular, help in understanding a model’s error rates. In the case of malware detection, it is particularly important to consider TPR, since a benign application labelled as malware is less likely to cause as many problems as malware labelled as benign.
@@
The most common approach to comparing ML models in the Android malware literature is to report the average of metrics after using a resampling method like -fold cross validation. Following this approach, we use standard 10-fold cross-validation (CV), and report the averages for all metrics. However, the average of a distribution can be misleading, particularly when there is a high variance between folds. To account for this, we report standard deviations and also carry out statistical tests when comparing averages of different approaches for malware detection. In a statistical test, a null hypothesis is set up (H0): the average is equal for all sample population groups. A significant difference between the samples is found if the results from the test were to reject the null hypothesis and accept the alternative hypothesis (Ha): the average is not equal for all sample population groups.
@@
In this work, we use the Kruskal–Wallis (Kruskal and Wallis, 1952) omnibus test followed by pairwise Dunn’s (Dunn, 1964) tests. Kruskal–Wallis is a non-parametric test used to determine if there are significant differences between the medians of two or more distributions. In this case, the distributions are the set of metric scores produced by each ML model across folds. If the test statistic p-value is less than a threshold level, the null hypothesis is rejected, indicating that there is a significant difference between the groups. Kruskal-Wallis is an omnibus test and so does not indicate which sample group is statistically different from another. If Kruskal–Wallis finds a significant difference, Dunn’s test is then used to identify the samples that are statistically different from the others. If more than two models are being compared, a pairwise tournament between the ML models is used to determine the ML models with significant statistical differences. Dunn’s test automatically adjusts for multiple comparisons when used with Kruskal-Wallis. We use a threshold of p=0.05 for all statistical tests, i.e. a confidence of 95%.
2.4. Choice of Features and Classes
We used the tool described in Section 2.2 to extract features for Android malware detection. These features were chosen following a comprehensive review of the literature, as outlined in (Muzaffar et al., 2022), focusing on features which have been widely used in previous studies. We also took into account the findings of other prominent review papers (Naway and LI, 2018; Alqahtani et al., 2019; Bayazit et al., 2020; Wang et al., 2020).
In this paper, we focus on the binary classification task of discriminating malware from benign applications. However, some studies (e.g. Son et al. (2021, 2022)) consider multi-class problems, typically with the aim of discriminating between different malware families. We choose to focus on binary classification for several reasons. First, the aim of this study is to understand the role of different models and features in the malware detection process, and binary classification provides the most generalisable insights into this. Second, there is no universally-accepted method for labeling malware families, with the ever-changing and diverse nature of malware leading to varying taxonomies and classifications across the cybersecurity community. Again, this poses difficulties when trying to reach generalisable insights. Third, there is a significant imbalance within the natural distribution of malware types. For instance, in the dataset collected for this study, adware comprises around 70% of the samples. This introduces a challenge of handling imbalanced data, but also the challenge of separating the effects of mechanisms used to address imbalance from the effects of models and features.
2.5. Core ML Models and Feature Selection Algorithms
Previous studies have used a diverse range of ML models. In general, we reimplement the model(s) used in the original study wherever possible. However, to provide better consistency and comparability across studies, we also augment the original studies with the following core set of ML models, if they are not already included: support vector machine (SVM), random forest (RF), decision tree (DT) and -nearest neighbour (kNN) (Hastie et al., 2009). All of these are widely used both within the Android malware detection literature (as surveyed in (Muzaffar et al., 2022)) and the broader ML literature. We use grid search to tune the hyperparameters of all the models implemented. This is to ensure fairness when making comparisons between models (Lones, 2021).
3. Static Analysis
@@
3.1. Permissions and API calls
To begin with, we revisit one of the earliest studies that looked at permissions and API calls, carried out by Peiravian and Zhu (Peiravian and Zhu, 2013) in 2013. The authors compared models built using permissions alone, API calls alone, and a combination of both. In each case, they trained an SVM, a decision tree and a bagging classifier. They did not mention the kernel used for the SVM, nor the base classifier used for bagging, so in our reimplementation we used the default choices of a linear kernel for SVM and decision trees as the base classifier. At the time of their study, the total number of permissions in Android was 130 and the total number of API calls was only 1,326. As shown in Table 1, the numbers of both have grown since then, and by two orders of magnitude in the case of API calls.
@@ 3.1. Permissions and API calls
3.2. A Closer Look at Permissions
Nevertheless, permissions are easier to extract than other static features. For this reason, they are widely used in practice. To provide more insight into the best way to use permissions for Android malware detection, we reimplemented a study by Wang et al. (Wang et al., 2014) that aimed to provide a more fine-grained view of the risk associated with permissions. Their study used two groups of Android permissions to build ML models: normal permissions, which are automatically granted by the system, and dangerous permissions, which provide access to private data or allow control over the device. For more information about these categories, see (Android Developers, 2024b). As shown in Table 3, the total number of normal and dangerous permissions is only slightly different in the current Android release.
@@ 3.2. A Closer Look at Permissions
Rathore et al. (Rathore et al., 2021) further experimented with different feature selection and feature dimensionality reduction methods, with the aim of building an efficient model that used a minimal set of permissions. Apart from the original dataset with full permissions (OD), they used VT, PCA, and autoencoders (with one and three layers: AE-1L,AE-3L) to construct reduced feature sets. To this, we added our core set of feature selection algorithms. Unlike Wang et al., they carried out feature selection on the entire set of Android permissions, not just the normal and dangerous permissions. They trained six ML models for each of the feature selection/reduction methods: DT, kNN, SVM, RF, AdaBoost and DNNs (with two, four and seven layers). These were compared against models trained on the full set of 166 permissions.
@@
We reimplemented one more study that used only permissions. This study, by Sahin et al. (Şahin et al., 2023), is notable for using model-based feature selection. Unlike the more traditional forms of feature selection used in the previous studies, model-based feature selection uses an ML model to determine the importance of each feature. This can lead to more appropriate feature combinations; however, it does have a cost, and Sahin et al. attempt to mitigate against this by using a relatively inexpensive model, namely multiple regression, to assess features.
@@
| Model | Permissions | Accuracy | Precision | F1-Score | TPR | TNR | Permissions | Accuracy | Precision | F1-Score | TPR | TNR |
@@
To conclude, reasonable levels of performance can be achieved with permissions-based approaches, especially when the number of permissions is reduced using a feature selection algorithm, and RFs are used as the ML model. However, none of the permissions-based models we looked at were competitive against API call-based models.
3.3. Representations of API Calls
Next, we take a closer look at the best way of building models based on API calls, beginning with a reimplementation of the research carried out by Ma et al. (Ma et al., 2019), who considered three different ways of representing feature usage within an application: API usage, API frequency and API sequences.
For API usage, each API call is represented by a binary feature indicating whether the call is used within an application. We created our version of this dataset by using the complete set of 134,207 API calls; i.e., each application is represented by a feature vector , making this a high-dimensional feature set. This is the same dataset used for our earlier reimplementation of Peiravian et al. (Peiravian and Zhu, 2013). The API frequency data set is similar, except that each feature has an integer value, and indicates how many times each call is made within an application. For example, the feature vector for an application would be if and the application used 5 times, 4 times and 0 times.
@@ 3.3. Representations of API Calls
| Dataset | Model | Accuracy | Precision | F1-Score | TPR | TNR | Model | Accuracy | Precision | F1-Score | TPR | TNR |
| Usage | DT | - | 0.968 | 0.965 | 0.962 | - | Random Forest | 0.966 | 0.968 | 0.966 | 0.964 | 0.969 |
@@
| Ensemble | Voting | - | 0.991 | 0.99 | 0.988 | - | Voting | 0.965 | 0.963 | 0.962 | 0.967 | 0.963 |
@@
Ma et al. (Ma et al., 2019) used a different model for the three datasets: a DT for the API usage dataset, a deep neural network (DNN) for the API frequency dataset, and a long short-term memory (LSTM) model for the API sequence dataset. They also built an ensemble from the three models. In Table 7, we summarise our results from using this same approach. This indicates that the API frequency model performs the best, with both the usage and frequency models outperforming the API sequence models. We did not find any benefit to ensembling the models. This is in opposition to the findings of Ma et al. (Ma et al., 2019), for whom the best model was the ensemble, followed by the sequence model. Like Ma et al., we also looked more closely at the effect of network depth on the performance of the DNN and LSTM models. Table 8 shows the results, showing that best performance is now achieved with a DNN of 16 layers, which is twice the depth of the best DNN found by Ma et al.
@@
A limitation of the original study was that different models were used for different representations, meaning that the effect of changing the representation could not be isolated from the effect of changing the model. To give more insight into whether the additional information embedded in the frequency representation over the usage representation is useful, we trained each of our core models, plus DNN, on both the API usage and frequency datasets. The results are shown in Tables 11 and 9 respectively. Two things are notable: first, the differences in model accuracies are relatively small. Second, the ranking of the models changes depending on the representation, with DNNs doing best on the frequency dataset and RFs doing best on the usage dataset. Perhaps coincidentally, this justifies the model choices of Ma et al.
3.4. Reducing the Number of API Calls
@@
Jung et al. (Jung et al., 2018) approached this by analysing the top 50 API calls used in benign applications then malware applications, resulting in two feature sets, one from each of the applications type, which they used to train RF models. The approach was evaluated using a dataset of 30,159 benign and 30,084 malicious applications. The results of our reimplementation are shown in Table 10. Of the two feature sets, the use of the top 50 malware API calls leads to better models, and this is supported by a Kruskal-Wallis test, which showed a significant difference (H = 14.23, p 0.05). However, the rates of accuracy achieved in our experiments are considerably less than those reported by Jung et al, and the resulting models performed relatively poorly. This difference may in part be due to the much larger size of the API call set in contemporary versions of Android.
@@
| Features | Features | Accuracy | Precision | F1-Score | TPR | TNR | Features | Accuracy | Precision | F1-Score | TPR | TNR |
@@
| Chi-Square | - | - | - | - | - | - | 10,000 | 0.961 | 0.961 | 0.954 | 0.967 | 0.953 |
Although these results suggest that 50 API calls are insufficient to train accurate models, from a practical perspective, it remains desirable to work with smaller feature sets. Consequently, in Muzaffar et al. (Muzaffar et al., 2021), we investigated the use of dimensionality reduction in API feature sets more broadly. We used a dataset of 20,000 benign applications and 20,000 malicious applications, and our results showed that the accuracy of the ML models could be improved by reducing the number of API calls by approximately 95% from the original set of 134,207. We trained SVM, RF, DT, AdaBoost and Naïve Bayes models, and used MI, VT and PCC for feature selection.
@@
Regardless of feature set size, RFs were the best models. Table 12 shows the RF performance metrics for each of the feature selection algorithms. A Kruskal-Wallis test between the best performing RF model trained on the full API set and the models from Table 12 indicated significant differences (H = 32.59, p 0.05). Post hoc Dunn’s tests then showed that feature sets derived using variance threshold led to significantly better models. Notably, these models could outperform the full API set models across all metrics while using only 5% (6,443) of the total API calls.
@@
3.5. The Drebin feature set
The Drebin dataset of Arp et al. (Arp et al., 2014) is widely used in Android malware detection, both due to its availability and the large number of extracted features. They provided the following features for each application in the dataset:
@@ 3.5. The Drebin feature set
Used permissions: Felt et al. (Felt et al., 2011) introduced a method which was used to match API calls to permissions, hence obtaining the permissions which are actually used by the application.
@@
| Benign | Malware | No. of features | Accuracy | |
@@
3.6. Opcodes
Rather than a sequence of API calls, a program can instead be viewed, at a lower level, as a sequence of machine language instructions, or opcodes. A relatively small number of studies have considered opcode analysis as a basis for Android malware detection, and most have involved extracting features in the form of n-grams of opcodes, referred to as n-opcodes. For example, lets say that an application is the sequence , , , . A 2-gram, also known as bigram, representation of would be three bigrams, namely [, ], [, ] and [, ]. To measure the benefits of this approach, and compare it against more traditional static features like permissions and API calls, we reimplement a study by Kang et al. (Kang et al., 2016), which considered both usage and frequency representations (see Section 3.3) of n-opcodes.
@@ 3.6. Opcodes
| Kang et al. | Our Dataset | ||
| n | n-opcodes | n | n-opcodes |
| 1 | 214 | 1 | 218 |
| 2 | 22,371 | 2 | 44,411 |
| 3 | 399,598 | 3 | 1,049,741 |
| 4 | 2,201,377 | 4 | 8,474,254 |
| 5 | 6,458,246 | 5 | 35,033,022 |
| 6 | 12,969,857 | 6 | 95,082,086 |
| 7 | 20,404,473 | 7 | 195,415,174 |
| 8 | 27,366,890 | 8 | 330,418,164 |
| 9 | 33,024,116 | 9 | 484,711,644 |
@@
Table 17 summarises the performance of the best models. A Kruskal-Wallis test applied to these results indicates a significant difference (H = 36.44, p 0.05) between models, with Dunn’s tests showing that RF and SVM models are significantly better than the others. The performance of our models was broadly similar to those reported by Kang et al. Notably, these models have similar performance to models based on API calls. Opcodes take less time to extract compared to API calls and are much easier to train computationally on the base feature sets. However, it is much easier to interpret the API call results as APIs correspond to specific features in the Android SDK.
| Kang et al. | Our Results | ||||||||||||
| Classifier | n | Accuracy | Precision | F1-Score | TPR | TNR | Classifier | n | Accuracy | Precision | F1-Score | TPR | TNR |
| Usage | Usage | ||||||||||||
| SVM | 3 | - | - | 0.98 | - | - | SVM | 4 | 0.968 | 0.971 | 0.971 | 0.961 | 0.971 |
| Random Forest | 4 | - | - | 0.98 | - | - | Random Forest | 5 | 0.973 | 0.979 | 0.975 | 0.965 | 0.979 |
| Decision Trees | 5 | - | - | 0.98 | - | - | Decision Trees | 5 | 0.952 | 0.952 | 0.951 | 0.949 | 0.952 |
| Naïve Bayes | 3 | - | - | 0.84 | - | - | Naïve Bayes | 5 | 0.832 | 0.834 | 0.828 | 0.821 | 0.838 |
| Frequency | Frequency | ||||||||||||
| SVM | 4 | - | - | 0.96 | - | - | SVM | 5 | 0.956 | 0.954 | 0.956 | 0.958 | 0.954 |
| Random Forest | 8 | - | - | 0.97 | - | - | Random Forest | 5 | 0.965 | 0.969 | 0.967 | 0.961 | 0.970 |
| Decision Trees | 3 | - | - | 0.97 | - | - | Decision Trees | 6 | 0.940 | 0.942 | 0.937 | 0.935 | 0.941 |
| Naïve Bayes | 4 | - | - | 0.85 | - | - | Naïve Bayes | 5 | 0.791 | 0.794 | 0.784 | 0.781 | 0.798 |
@@
A common approach within deep learning is to convert numerical data into pseudoimages and use these to train a convolutional neural network (CNN). An example of this approach within the Android malware detection literature can be found in Xiao and Yang (Xiao and Yang, 2019), who converted disassembled opcode files into RGB images. To determine whether this has any advantage over n-opcode encodings, we reimplemented the same approach and wrote an algorithm that converts opcode sequence files into RGB images. We then used these to train a CNN model, using the same CNN topology as Xiao and Yang, which comprises standard convolutional, pooling and fully-connected layers. Xiao and Yang trained and evaluated their model using a dataset of 4,406 benign and 6,134 malicious applications. Table 18 compares their results against those of our reimplementation, which reported slightly lower metric scores. These shows that there is generally no benefit to using this more complex modelling approach over the simpler models reported in Table 17. This is supported by a statistical analysis, with Kruskal-Wallis (H = 56.44, p 0.05) and Dunn’s test showing Xiao and Yang’s approach to be significantly worse on our dataset than the best models in Table 17.
@@
Another example of using deep learning for opcode analysis can be found in a study by Yeboah and Baz Musah (Yeboah and Baz Musah, 2022). In this case, they used a 1D CNN; that is, a one-dimensional CNN which can be applied to sequential numeric data. A notable element of their approach was the use of a word2vec embedding model to represent each opcode as a single numeric value, allowing opcode sequences to be represented as dense real-valued vectors (as opposed to using one-hot encodings). The embedding model was trained separately on unigram and bigram opcode sequences, resulting in matrices of size M×N, where M is the number of unique opcodes and N is the vector dimension. N was set to 64 for both unigram and bigram embeddings. The 1D CNN comprised convolutional layers, pooling layers, and fully connected layers. Table 19 shows the results of our reimplementation, and again shows that there is no benefit to using a more complex modelling approach over the simpler models reported in Table 17. Our results were significantly lower than those reported by Yeboah and Baz Musah, who used a small (4948 malware, 2477 benign) though relatively recent dataset.

### v4

@@ 4.5 Model evaluation procedure
(2) - True Positive Rate (TPR)
@@
(4) - True Negative Rate (TNR)
@@
In general, it is necessary to use a range of metrics in order to give a full picture of the performance of different ML models [Lones2021]. With this in mind, we report the accuracy, precision, F1-score, TPR and TNR for each model. TPR and TNR, in particular, help in understanding a model’s error rates. In the case of malware detection, it is particularly important to consider TPR, since a benign application labelled as malware is less likely to cause as many problems as malware labelled as benign.
The most common approach to comparing ML models in the Android malware literature is to report the average of metrics after using a resampling method like -fold cross validation. Following this approach, we use standard 10-fold cross-validation (CV), and report the averages for all metrics. The choice of 10 folds gives a good balance between bias and variance. It is computationally efficient, especially when dealing with complex feature sets, and is widely adopted in similar malware detection studies, ensuring reliable and reproducible results.
@@
However, the average of a distribution can be misleading, particularly when there is a high variance between folds. To account for this, we report standard deviations and also carry out statistical tests when comparing averages of different approaches for malware detection. In a statistical test, a null hypothesis is set up (H0): the average is equal for all sample population groups. A significant difference between the samples is found if the results from the test were to reject the null hypothesis and accept the alternative hypothesis (Ha): the average is not equal for all sample population groups.
@@
In this work, we use the Kruskal–Wallis [Kruskal1952] omnibus test followed by pairwise Dunn’s [Dunn1964] tests. Kruskal–Wallis is used to determine if there are significant differences between the medians of two or more distributions. In this case, the distributions are the set of metric scores produced by each ML model across folds. Kruskal–Wallis is non-parametric, so does not assume that the data fits a normal distribution. This is important because the distribution of accuracies from k-fold CV may not be normally distributed, and there is no reliable means of measuring normality in small samples. If the test statistic p-value is less than a threshold level, the null hypothesis is rejected, indicating that there is a significant difference between the groups. Kruskal-Wallis is an omnibus test and so does not indicate which sample group is statistically different from another. If Kruskal–Wallis finds a significant difference at the group level, Dunn’s test is then used to identify the samples that are statistically different from the others. Dunn’s is a pairwise test, i.e. for the CV distributions of two models, it determines whether there is a significant difference in their medians. If there are more than two models in a group, a pairwise tournament between all ML models is used to those which exhibit significant statistical differences from the others. Dunn’s test automatically adjusts for multiple comparisons when used with Kruskal-Wallis. We use a threshold of p=0.05 for all statistical tests, i.e. a confidence of 95%. The previous studies we review did not, in general, carry out statistical tests, but these are important in order to reach reliable conclusions.
5 Reimplementation and extension of static modelling approaches
@@
5.1 Use of permissions and API call features
To begin with, we revisit one of the earliest studies that looked at permissions and API calls, carried out by Peiravian and Zhu [Peiravian2013] in 2013. The authors compared models built using permissions alone, API calls alone, and a combination of both. In each case, they trained an SVM, a decision tree and a bagging classifier. They did not mention the kernel used for the SVM, nor the base classifier used for bagging, so in our reimplementation we used the default choices of a linear kernel for SVM and decision trees as the base classifier. At the time of their study, the total number of permissions in Android was 130 and the total number of API calls was only 1,326. As shown in Table 1, the numbers of both have grown since then, and by two orders of magnitude in the case of API calls.
@@ 5.1 Use of permissions and API call features
5.2 Selection of permissions feature sets
Nevertheless, permissions are easier to extract than other static features. For this reason, they are widely used in practice. To provide more insight into the best way to use permissions for Android malware detection, we reimplemented a study by Wang et al. [Wang2014] that aimed to provide a more fine-grained view of the risk associated with permissions. Their study used two groups of Android permissions to build ML models: normal permissions, which are automatically granted by the system, and dangerous permissions, which provide access to private data or allow control over the device. For more information about these categories, see [android_permission]. Table 3 shows the number of normal and dangerous permissions as tagged by Android compared to the full permission set as shown in Table 1.
@@ 5.2 Selection of permissions feature sets
Rathore et al. [Rathore2021] further experimented with different feature selection and feature dimensionality reduction methods, with the aim of building an efficient model that used a minimal set of permissions. Apart from the original dataset with full permissions (OD), they used VT, PCA, and autoencoders (with one and three layers: AE-1L,AE-3L) to construct reduced feature sets. To this, we added our core set of feature selection algorithms. Unlike Wang et al., they carried out feature selection on the entire set of Android permissions, not just the normal and dangerous permissions. They trained six ML models for each of the feature selection/reduction methods: DT, kNN, SVM, RF, AdaBoost and DNNs (with two, four and seven layers). These were compared against models trained on the full set of 166 permissions.
@@
We reimplemented one more study that used only permissions. This study, by Sahin et al. [Sahin2023], is notable for using model-based feature selection. Unlike the more traditional forms of feature selection used in the previous studies, model-based feature selection uses an ML model to determine the importance of each feature. This can lead to more appropriate feature combinations; however, it does have a cost, and Sahin et al. attempt to mitigate against this by using a relatively inexpensive model, namely multiple regression, to assess features.
@@
| Classifier | Permissions | Accuracy | Precision | F1-Score | TPR | TNR | Permissions | Accuracy | Precision | F1-Score | TPR | TNR |
@@
To conclude, reasonable levels of performance can be achieved with permissions-based approaches (see tables 4, 5 and 6), especially when the number of permissions is reduced using a feature selection algorithm, and RFs are used as the ML model. However, none of the permissions-based models we looked at were competitive against API call-based models (see table 2, and also the results in Section 5.3), with statistical analysis (Kruskal–Wallis H = 50.04, p 0.05) confirming that these differences are significant.
5.3 Comparison of API call representations
Next, we take a closer look at the best way of building models based on API calls, beginning with a reimplementation of the research carried out by Ma et al. [Ma2019], who considered three different ways of representing feature usage within an application: API usage, API frequency and API sequences.
For API usage, each API call is represented by a binary feature indicating whether the call is used within an application. We created our version of this dataset by using the complete set of 134,207 API calls; i.e., each application is represented by a feature vector , making this a high-dimensional feature set. This is the same dataset used for our earlier reimplementation of Peiravian et al. [Peiravian2013]. The API frequency data set is similar, except that each feature has an integer value, and indicates how many times each call is made within an application. For example, the feature vector for an application would be if and the application used 5 times, 4 times and 0 times.
@@ 5.3 Comparison of API call representations
| Usage | DT | - | 0.968 | 0.965 | 0.962 | - | RF | 0.966 | 0.968 | 0.966 | 0.964 | 0.969 |
@@
| Ensemble | Voting | - | 0.991 | 0.990 | 0.988 | - | Voting | 0.965 | 0.963 | 0.962 | 0.967 | 0.963 |
@@
Ma et al. [Ma2019] used a different model for the three datasets: a DT for the API usage dataset, a deep neural network (DNN) for the API frequency dataset, and a long short-term memory (LSTM) model for the API sequence dataset. They also built an ensemble from the three models. In Table 7, we summarise our results from using this same approach. This indicates that the API frequency model performs the best, with both the usage and frequency models outperforming the API sequence models. We did not find any benefit to ensembling the models. This is in opposition to the findings of Ma et al. [Ma2019], for whom the best model was the ensemble, followed by the sequence model. Like Ma et al., we also looked more closely at the effect of network depth on the performance of the DNN and LSTM models. Table 8 shows the results, showing that best performance is now achieved with a DNN of 16 layers, which is twice the depth of the best DNN found by Ma et al.
@@
A limitation of the original study was that different models were used for different representations, meaning that the effect of changing the representation could not be isolated from the effect of changing the model. To give more insight into whether the additional information embedded in the frequency representation over the usage representation is useful, we trained each of our core models, plus DNN, on both the API usage and frequency datasets. The results are shown in Table 9. Two things are notable: first, the differences in model accuracies are relatively small. Second, the ranking of the models changes depending on the representation, with DNNs doing best on the frequency dataset and RFs doing best on the usage dataset. Perhaps coincidentally, this justifies the model choices of Ma et al.
5.4 Selection of API call feature sets
@@
Jung et al. [Jung2018] approached this by analysing the top 50 API calls used in benign applications and malware applications, resulting in two feature sets, one from each of the applications type, which they used to train RF models. The approach was evaluated using a dataset of 30,159 benign and 30,084 malicious applications. The results of our reimplementation are shown in Table 10. Of the two feature sets, the use of the top 50 malware API calls leads to better models, and this is supported by a Kruskal-Wallis test, which showed a significant difference (H = 14.23, p 0.05). However, the rates of accuracy achieved in our experiments are considerably less than those reported by Jung et al, and the resulting models performed relatively poorly. This difference may in part be due to the much larger size of the API call set in contemporary versions of Android.
@@
| Selection | Features | Accuracy | Precision | F1-Score | TPR | TNR | Features | Accuracy | Precision | F1-Score | TPR | TNR |
@@
| CS | - | - | - | - | - | - | 10,000 | 0.961 | 0.961 | 0.954 | 0.967 | 0.953 |
| None | 134,207 | 0.959 | 0.960 | 0.959 | 0.957 | 0.960 | 134,207 | 0.966 | 0.968 | 0.966 | 0.964 | 0.969 |
Although these results suggest that 50 API calls are insufficient to train accurate models, from a practical perspective, it remains desirable to work with smaller feature sets. Consequently, in Muzaffar et al. [Muzaffar2021], we investigated the use of dimensionality reduction in API feature sets more broadly. We used a dataset of 20,000 benign applications and 20,000 malicious applications, and our results showed that the accuracy of the ML models could be improved by reducing the number of API calls by approximately 95% from the original set of 134,207. We trained SVM, RF, DT, AdaBoost and Naïve Bayes models, and used MI, VT and PCC for feature selection.
@@
Regardless of feature set size, RFs were the best models. Table 12 shows the RF performance metrics for each of the feature selection algorithms. A Kruskal-Wallis test across these models and the best performing RF model trained on the full API set indicated significant differences (H = 32.59, p 0.05). Post hoc Dunn’s tests then showed that feature sets derived using variance threshold led to significantly better models. Notably, these models could outperform the full API set models across all metrics while using only 5% (6,443) of the total API calls.
@@
5.5 Use of the Drebin feature set
The Drebin dataset of Arp et al. [Arp2014] is widely used in Android malware detection, both due to its availability and the large number of extracted features. They provided the following features for each application in the dataset:
@@ 5.5 Use of the Drebin feature set
Used permissions: Felt et al. [Felt] introduced a method which was used to match API calls to permissions, hence obtaining the permissions which are actually used by the application.
@@
| Benign | Malware | Number of features | Accuracy | |
@@
5.6 Use and representation of opcode features
Rather than a sequence of API calls, a program can instead be viewed, at a lower level, as a sequence of machine language instructions, or opcodes. A relatively small number of studies have considered opcode analysis as a basis for Android malware detection, and most have involved extracting features in the form of n-grams of opcodes, referred to as n-opcodes. For example, lets say that an application is the sequence , , , . A 2-gram, also known as bigram, representation of would be three bigrams, namely [, ], [, ] and [, ]. To measure the benefits of this approach, and compare it against more traditional static features like permissions and API calls, we reimplement a study by Kang et al. [Kang2016], which considered both usage and frequency representations (see Section 5.3) of n-opcodes.
@@ 5.6 Use and representation of opcode features
| n | Kang et al. | Our Dataset |
| 1 | 214 | 218 |
| 2 | 22,371 | 44,411 |
| 3 | 399,598 | 1,049,741 |
| 4 | 2,201,377 | 8,474,254 |
| 5 | 6,458,246 | 35,033,022 |
| 6 | 12,969,857 | 95,082,086 |
| 7 | 20,404,473 | 195,415,174 |
| 8 | 27,366,890 | 330,418,164 |
| 9 | 33,024,116 | 484,711,644 |
@@
Table 17 summarises the performance of the best models. The best value of across both representations and across most of the ML models is . This is a more consistent observation than in Kang et al’s original study, where the best value of varied between 3 and 8, with a preference towards 3 and 4. Given that the F1 scores from our reimplementation are broadly similar to those originally published, this may indicate a change in the underlying pattern of opcode usage within Android malware. However, the original dataset was relatively small, so there may also have been a degree of sample bias.
@@
A Kruskal-Wallis test applied to our results indicates a significant difference (H = 36.44, p 0.05) between models, with Dunn’s tests showing that RF and SVM models are significantly better than the others. Notably, these models have similar performance to models based on API calls. Opcodes take less time to extract compared to API calls and are much easier to train computationally on the base feature sets. However, it is much easier to interpret the API call results as APIs correspond to specific features in the Android SDK.
| Kang et al. | Our Results | |||||||||||
| Classifier | n | Accuracy | Precision | F1-Score | TPR | TNR | n | Accuracy | Precision | F1-Score | TPR | TNR |
| Features: Opcode usage | ||||||||||||
| SVM | 3 | - | - | 0.98 | - | - | 4 | 0.968 0.002 | 0.971 0.001 | 0.971 0.002 | 0.961 0.001 | 0.971 0.002 |
| RF | 4 | - | - | 0.98 | - | - | 5 | 0.973 0.002 | 0.979 0.002 | 0.975 0.001 | 0.965 0.001 | 0.979 0.002 |
| DT | 5 | - | - | 0.98 | - | - | 5 | 0.952 0.001 | 0.952 0.001 | 0.951 0.003 | 0.949 0.002 | 0.952 0.001 |
| Naïve Bayes | 3 | - | - | 0.84 | - | - | 5 | 0.832 0.002 | 0.834 0.002 | 0.828 0.001 | 0.821 0.003 | 0.838 0.002 |
| Features: Opcode frequency | ||||||||||||
| SVM | 4 | - | - | 0.96 | - | - | 5 | 0.956 0.001 | 0.954 0.002 | 0.956 0.002 | 0.958 0.003 | 0.954 0.003 |
| RF | 8 | - | - | 0.97 | - | - | 5 | 0.965 0.002 | 0.969 0.003 | 0.967 0.002 | 0.961 0.002 | 0.970 0.002 |
| DT | 3 | - | - | 0.97 | - | - | 6 | 0.940 0.001 | 0.942 0.001 | 0.937 0.002 | 0.935 0.003 | 0.941 0.001 |
| Naïve Bayes | 4 | - | - | 0.85 | - | - | 5 | 0.791 0.002 | 0.794 0.001 | 0.784 0.003 | 0.781 0.002 | 0.798 0.001 |
@@
A common approach within deep learning is to convert numerical data into pseudoimages and use these to train a convolutional neural network (CNN). An example of this approach within the Android malware detection literature can be found in Xiao and Yang [Xiao2019a], who converted disassembled opcode files into RGB images. To determine whether this has any advantage over n-opcode encodings, we reimplemented the same approach and wrote an algorithm that converts opcode sequence files into RGB images. We then used these to train a CNN model, using the same CNN topology as Xiao and Yang, which comprises standard convolutional, pooling and fully-connected layers. Xiao and Yang trained and evaluated their model using a dataset of 4,406 benign and 6,134 malicious applications. Table 18 compares their results against those of our reimplementation, which reported slightly lower metric scores. This deviation may be due to the restricted focus of the original study, which only considered three malware families due to data sparsity.
Compared to Table 17, the results of our reimplementation are lower across the board than when using Kang et al’s n-opcode representation. This shows that there is generally no benefit to using this more complex modelling approach over simpler models. Rather, it appears that either the direct sequential encoding of opcodes, the mapping to a pseudoimage, or the particular nature of the CNN model, makes the underlying discriminative pattern more difficult to detect. This observation is supported by a statistical analysis, with Kruskal-Wallis (H = 56.44, p 0.05) and Dunn’s test showing Xiao and Yang’s approach to be significantly worse on our dataset than the best models in Table 17.
@@
Another example of using deep learning for opcode analysis can be found in a study by Yeboah and Baz Musah [yeboah2022]. In this case, they used a 1D CNN; that is, a one-dimensional CNN which can be applied to sequential numeric data. A notable element of their approach was the use of a word2vec embedding model to represent each opcode as a single numeric value, allowing opcode sequences to be represented as dense real-valued vectors (as opposed to using one-hot encodings). The embedding model was trained separately on unigram and bigram opcode sequences, resulting in matrices of size M×N, where M is the number of unique opcodes and N is the vector dimension. N was set to 64 for both unigram and bigram embeddings. The 1D CNN comprised convolutional layers, pooling layers, and fully connected layers. Table 19 shows the results of our reimplementation, and again shows that there is no benefit to using a more complex modelling approach over the simpler models reported in Table 17. Our results were significantly lower than those reported by Yeboah and Baz Musah, who used a small (4948 malware, 2477 benign) though relatively recent dataset.